If all the water were allowed to evaporate from 12 mL of a 0.64 mol/L solution of potassium iodide, KI, what mass of solute would remain?

Question 10 options:

106.24 g


1.99 g


12.75 g


1.27 g
1
Expert's answer
2020-07-20T06:26:09-0400

 12 mL of a 0.64 mol/L solution of potassium iodide Have


number of moles of KI ( Solute ) = ( 12 * 0.64 ) / 1000 . = 0.00768 MOLE .


MOLAR MASS OF kI = 127 + 39 = 166 g / mole .


So , Mass of KI = ( 0.00768 * 166 ) g . = " 1.27 g" ANSWER ......